# Estimating LLM Consistency: A User Baseline vs Surrogate Metrics

### v2

@@ Abstract
Large language models (LLMs) are prone to hallucinations and sensitive to prompt perturbations, often resulting in inconsistent or unreliable generated text. Different methods have been proposed to mitigate such hallucinations and fragility—one of them being measuring the consistency (the model’s confidence in the response, or likelihood of generating a similar response when resampled) of LLM responses. In previous work, measuring consistency often relied on the probability of a response appearing within a pool of resampled responses, or internal states or logits of responses. However, it is not yet clear how well these approaches approximate how humans perceive the consistency of LLM responses. We performed a user study ($n=2,976$) and found current methods typically do not approximate users’ perceptions of LLM consistency very well. We propose a logit-based ensemble method for estimating LLM consistency, and we show that this method matches the performance of the best-performing existing metric in estimating human ratings of LLM consistency. Our results suggest that methods of estimating LLM consistency without human evaluation are sufficiently imperfect that we suggest evaluation with human input be more broadly used.
@@
Xiaoyuan Wu1. Weiran Lin1, Omer Akgul2, 1, Lujo Bauer1 1Carnegie Mellon University 2RSAC Labs Correspondence:wxyowen@cmu.edu
@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption in a multitude of domains despite numerous inherent limitations such as hallucinations and fragility against adversarial attacks. Hallucinations can have disastrous consequences in high-stakes fields like healthcare and legal Merken (2025), prompting concern even as adoption continues Tiffin and Fraser (2025); Metnick et al. (2024). Further, the fragility of models enables a range of misuses that could harm users Kumar and Lakkaraju (2024); Lin et al. (2025). For instance, Lin et al. (2025) show that imperceptible changes to suggested prompts can result in outputs with biases controlled by an adversary.
@@
Researchers have suggested that some of these issues correlate with the inconsistency of the LLMs, which is generally defined as the tendency to generate low-confidence responses or conflicting responses when the same prompt is resampled. An accurate estimation of consistency can be used: to predict if a given answer is factual, enabling the higher accuracy needed in high-stakes domains Duan et al. (2024); to detect fragile (or perhaps malicious) prompts Lin et al. (2025); in membership-inference attacks Mattern et al. (2023); to signal to users how much they should trust LLM outputs Kapoor et al. (2024); Ruggieri and Pugnana (2025).
@@
Motivated by such utility, many have tried to define and measure the consistency of LLMs. This body of work can roughly be divided into two categories: (1) estimating consistency based on the LLMs’ internal states or logits and (2) estimating consistency based on resampling the LLMs’ responses. While the former is more compute efficient, the latter is empirically well-grounded Kuhn et al. (2023); Qiu and Miikkulainen (2024) and applicable even without white-box access to the model Lin et al. (2024).
@@
Sampling-based estimation methods fundamentally rely on sampling from an LLM-prompt pair and comparing the outputs using a comparison function. This function differs between works—e.g., Duan et al. (2024) vs Manakul et al. (2023)—but partially boils down to estimating semantic similarity of responses. However, to our knowledge, none have based their estimation of model consistency on user-based comparisons, the ground truth for semantic similarity Bowman et al. (2015); Agirre et al. (2014). Further, the plethora of proposed uncertainty estimation methods has not yielded a clear baseline consistency metric. Without a baseline, consistency metrics are typically evaluated by their ability to predict whether a model output is factual Kuhn et al. (2023); Qiu and Miikkulainen (2024); Duan et al. (2024); Zhang et al. (2024); Kapoor et al. (2024). Joining recent work Novikova et al. (2025), we argue that consistency might be able to predict accuracy but is fundamentally an independent property of the model-prompt pair. When fed a prompt, there should be a metric that conveys to what degree the model will produce responses with equivalent meaning. We further posit that the baseline for measuring consistency is reliant on the comparison of semantics, the ground truth of which is defined by human judgement Bowman et al. (2015); Agirre et al. (2014); Nguyen et al. (2014); Raj et al. (2025).
@@
We aim to fill this gap by using user-based comparisons to estimate the consistency of LLMs. Specifically, we conduct a user study with 2,976 participants to collect semantic similarity ratings between a sample of 10 responses to each of 100 prompts, totalling 14,880 comparisons. We then calculate response-level (one score per response) and prompt-level (one score per prompt) consistency, establishing an user-based LLM consistency baseline. Through a series of experiments, we show that existing metrics for uncertainty do not align well with human judgements collected in this study. We further show that an ensemble of logit-based scores is as similar to human judgement as the best-performing of the other methods we tested. We find that the discrepancy between existing metrics and human judgements fluctuates between models and between datasets, with previous metrics being less similar to human judgements on real-world prompts then on synthetic prompts. Based on our results we advocate for more human-based consistency evaluation in future work.
@@
We structure the remainder of the paper in four sections: 2 details background work; 3 describes our model- and prompt-selection strategy, user study, and consistency calculation; 4 reports our experiments and comparisons; finally, we summarize and discuss implications of our work in 5.
@@ 2 Related Work
Here we review work in consistency estimation of LLMs with resampling and internal model states. (2.1). We also cover work that investigated logit-based metrics to estimate LLM consistency (2.2). Finally, we reiterate our study motivation (2.3).
@@ 2.2 Probability- and Logit-Based Uncertainty Metrics
Existing work suggests the confidence of individual token generation can suggest the consistency of LLM responses Manakul et al. (2023). Borrowing definitions from Manakul et al. (2023), we denote the probability distribution of an individual token generated as $p$, and the entropy of individual token generation $H$, $H=-\sum_{i}p_{i}log(p_{i})$. Existing metrics include the average of the minus log probability $Avg(-log(p))$, maximum of the minus log probability $Max(-log(p))$, average of entropy $Avg(H)$ and maximum of entropy $Max(H)$. In this paper, we ensemble these metrics (and variants) in 3.3.
@@
Besides probability-based (i.e., $p$-based) uncertainty metrics, we also try to find logit-based uncertainty metrics to suggest LLM response consistency. We denote the logits of individual token generation as $l$, $p=softmax(l)$. Specifically, we denote the largest, second largest, third largest, and fourth largest logit of a token generation as $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ correspondingly. The Difference of Logits Ratio Loss, i.e., the DLR loss, formally $\frac{l_{1}-l_{2}}{[l_{1}-(l_{3}+l_{4})]/2}$, is a loss function commonly used by adversarial attacks to suggest how confident a model is Croce and Hein (2020). In this paper, we borrow the DLR loss as a logit-based uncertainty of a token generation to estimate the consistency of LLM responses.
Since probability-based uncertainty estimation is ultimately derived from token logits, for the rest of the paper we call this group of uncertainty estimation logit-based methods.
@@ 2.3 Importance of User Evaluation in NLP Tasks
Human evaluations are important in improving and assessing the quality of natural language processing (NLP) Gritta et al. (2024); Boyd-Graber et al. (2022); Blodgett et al. (2024). Researchers have investigated how well machines are able to translate sentences into another language, compare sentences for semantic similarity, and many other NLP tasks with human evaluations Chatzikoumi (2020); Graham et al. (2017). This trend has continued in the LLM era Ouyang et al. (2022); Bai et al. (2022) While prior work has found ways to estimate how consistent responses produced by LLMs are, we argue that consistency is defined by what degree humans see responses as similar to each other. Using this definition, we establish a baseline through a user study.
@@ 3 Methods
In this study, we seek to understand how good existing metrics can approximate users’ perceptions of LLM consistency. We explore ways to create novel and more efficient ways of estimating LLM consistency without the need for expensive user studies or generating multiple responses per prompt. We provide an overview of our study design in figure 1.
In this section we first describe the prompts, models, and responses used in the study (3.1). We then describe in detail how we conducted the user study with 2,976 participants (3.2). Finally, we explain how we calculate consistency, both with existing metrics and our user-based method (3.3).
@@ 3.1 Prompts, Models, and Responses
Studies of LLMs’ consistency fundamentally rely on a fixed set of prompts Geng et al. (2024). In this work, we selected two prompt sets: (1) To ensure comparability with prior research, we use a dataset of open-ended questions from CoQA Reddy et al. (2019), which is commonly adopted in the literature; and (2) To better represent real-world use cases, we include a sample of 50 prompts from LMSYS-Chat-1M Zheng et al. (2023).
Each entry in the CoQA dataset contains a story with a series of questions about the story. For our study, we only used the first question in each series as many following questions are dependent on the previous question or answer for context. We randomly sampled 50 story-question pairs from the CoQA dataset.
Real-world prompts from the LMSYS dataset are more diverse compared to those in CoQA. As such, we used labels in the data to filter out prompts that are non-English or Flagged111by OpenAI Moderation Zheng et al. (2023) (e.g., harassment, violence). Next, we randomly sampled 100 prompts from the filtered LMSYS dataset and manually removed 48 prompts – 14 coding questions, 9 about the LLM itself, 9 that were inappropriate (e.g., sexual), 6 that contain nonsensical sentences, 5 that we expected to have non-English responses (e.g., translations), 3 asking for answers over 1,500 words, and 2 that contain time sensitive information. We excluded coding questions since participants might not have the background to rate coding-related responses. We randomly removed 2 prompts from the remaining 52 to achieve a final sample of 50.
@@
A generalizable definition of consistency should be model-agnostic. Thus, we used three open-weights models from competeing institutions for generating responses: Llama-3.2-3B-Instruct (Llama), Gemma-2-9B-it (Gemma), and Mistral-7B-Instruct-v0.3 (Mistral). All three were the most recent publically available versions of their repective families, were widely used in prveious work, and met our hardware memory constraints (Nvidia RTX A6000). We used open-weights models as they provide us access to logits, which are necessary for our calcualtion described in section 3.3, and enables reproducibility of our study Ma et al. (2024). We left temperature settings unchanged from the default configuration because there is no consensus from pervious work on optimal temprature for uncertainty estimation Cecere et al. (2025); Du et al. (2025); Wang et al. (2023); Renze and Guven (2024); Zou et al. (2023)
@@
The 100 prompts were randomly assigned to the three LLMs, Gemma receiving 34, Llama 33, and Mistral 33. Using the assigned model, we then generated 10 responses per prompt, totaling 1,000 responses. We took this approach as previous work has shown 10 responses adequately captures the semantic diversity of the response space for uncertainty estimation purposes Kuhn et al. (2023); Qiu and Miikkulainen (2024). Response generation took 30 minutes of GPU time.
@@ 3.2 User Study
While many metrics have been created to approximate LLM consistency, users’ perceptions of how consistent LLMs are has been largely missing from the literature. Our user-rating based approach establishes such baseline. We recruited users from Prolific and required participants 18 years or older, reside in the United States, read and type in English fluently, and have at least 95% approval rate on the platform. We use Prolific over MTurk because recent work has shown superior data quality Tang et al. (2022); Moss and Litman (2018).
Prolific users who are eligible encountered the post of our study and would be able to participate. Each participant was given an introduction about the study before seeing the instructions and examples of how to rate semantic similarity between a pair of sentences using a 6-point scale (section A.1). This commonly used 6-point scale, with explanations and examples were adopted from Agirre et al. (2014), asking crowdsourced users to rate similarity between pairs of sentences. After the instructions, each participant rated the semantic similarity of five pairs of sentences from five different prompts. Lastly, we collected participants’ demographics. The sentence pairs and their orders within the survey were randomized. Our study was approved by our institution’s ethics review board.
@@
For each prompt, there are $\binom{num\_responses}{2}$ number of pairs of responses. For the 100 prompts, we obtained 1,000 responses, of which 756 are unique. These responses made up 4,500 pairs of responses, of which 2,938 are unique. More details about number of prompts, responses, and pairs of unique responses for each model are described in table 1.
@@ 3.3 Calculating and Comparing Consistency
Here we outline how we calculate consistency scores (we calculate similarities between pairs of responses, consistency for each response, and consistency for each prompt) and how we compare user-based scores to previous work.
@@
We first aggregate user ratings of similarity between a pair of responses $r_{a}$ and $r_{b}$ as the average ratings of $n\geq 5$ participants removing the highest and lowest scores Curran (2016) formally,
@@
To understand how existing metrics used for LLM consistency compare to human ratings, we compare them at the response-pair level. For each prompt (e.g., “What is Wh in batteries?”), 10 responses make up 45 response-pairs making a total of 4,500 response-pairs for the 100 prompts in our study. For these response-pairs, we calculate a Spearman correlation coefficient $\rho$ between $s_{\text{user}}$ and each of the four existing metrics—$s_{\text{\emph{Bert}}}$, $s_{\text{\emph{BLEU}}}$, $s_{\text{\emph{Rouge}}}$, and $s_{\text{\emph{USE}}}$. We use Spearman correlation coefficient $\rho$ because human ratings were on a 6-point likert scale, and we do not assume linear correlation between human ratings and existing metrics. We share our findings on how each of these compare to human ratings in section 4.2.
We calculated Bert using the BERTScore python package [^2] https://pypi.org/project/bert-score/0.3.0/, BLEU using NLTK python package [^3] https://www.nltk.org/_modules/nltk/translate/bleu_score.html, Rouge using the rouge-score python package [^4] https://pypi.org/project/rouge-score/, and USE using the universal-sentence-encoder model on Kaggle [^5] https://www.kaggle.com/models/google/universal-sentence-encoder.
@@
Given that $m$ responses are generated for one prompt, using similarity scores between pairs of responses, we calculate the response to response-set consistency by averaging the similarity between one response to all other $m-1$ responses to a prompt. Formally,
@@
In addition to the four metrics used in response-pair level comparison, we add Semantic Entropy (SEKuhn et al. (2023) at this level of comparison as their definition of consistency builds on the response to response-set level. To understand how well these five metrics perform at this level, we compute Spearman correlation coefficient and the mean squared error between each of the metrics and human ratings in section 4.3.
Additionally, motivated by reducing the overhead of resampling, we compare logits-based scores to human ratings and propose an ensembling method by using a linear combination of logits in section 4.3.
To calculate SE, we used the roberta-large-mnli model to check entailment between pairs of responses [^6] https://huggingface.co/FacebookAI/roberta-large-mnli. We provide more details on logit-based scores and our ensembling method in subsection 3.3.
@@
Building up from response to response-set consistency, we define prompt consistency as the average of the $m$ response to response-set consistency for all responses to the prompt, formally,
@@
We use Spearman correlation and mean squared error to determine how good each of the existing metrics and our ensemble is at approximating human ratings of prompt level consistency. We share our findings in section 4.4.
Estimating consistency with logit-based metrics.
We estimate the consistency of LLM responses using token-based uncertainty metrics, as previous work does Manakul et al. (2023). Specifically, we use four uncertainty metrics: the probability, minus log probability, entropy (all three introduced in 2.2), and the DLR loss (introduced in 2.2). For each response, we measure the maximum, sum, minimum, and average of these four token-based metrics, totaling $16$ values. The maximum and minimum suggest the ranges on individual tokens, while the average suggests the expected value across all tokens. The sum suggests a cumulative estimate over the whole response, as different responses may have different lengths. Unlike existing work that detects hallucination in one response Manakul et al. (2023), we collect several responses to estimate the consistency of LLM for a prompt, corresponding to the definition we give in equation 3.
@@
In addition to individual uncertainty-based metrics described in the previous paragraph, we further perform an ensemble of these to approximate human ratings of LLM’s consistency. We used a Sequential Feature Selector scikit-learn developers (2025) to determine the most important metrics before ensemble a linear combination to approximate human ratings.

### v3

@@ Abstract
Large language models (LLMs) are prone to hallucinations and sensitive to prompt perturbations, often resulting in inconsistent or unreliable generated text. Different methods have been proposed to mitigate such hallucinations and fragility, one of which is to measure the consistency of LLM responses—the model’s confidence in the response or likelihood of generating a similar response when resampled. In previous work, measuring LLM response consistency often relied on calculating the probability of a response appearing within a pool of resampled responses, analyzing internal states, or evaluating logits of resopnses. However, it was not clear how well these approaches approximated users’ perceptions of consistency of LLM responses. To find out, we performed a user study ($n=2,976$) demonstrating that current methods for measuring LLM response consistency typically do not align well with humans’ perceptions of LLM consistency. We propose a logit-based ensemble method for estimating LLM consistency and show that our method matches the performance of the best-performing existing metric in estimating human ratings of LLM consistency. Our results suggest that methods for estimating LLM consistency without human evaluation are sufficiently imperfect to warrant broader use of evaluation with human input; this would avoid misjudging the adequacy of models because of the imperfections of automated consistency metrics.
@@
Xiaoyuan Wu1, Weiran Lin1, Omer Akgul2, 1, Lujo Bauer1 1Carnegie Mellon University 2RSAC Labs Correspondence: wxyowen@cmu.edu
@@ 1 Introduction
Large language models (LLMs) have seen rapid adoption across a multitude of domains despite numerous inherent limitations, such as hallucinating responses and being fragile to adversarial inputs. Hallucinations can have disastrous consequences in high-stakes fields like healthcare and law Merken (2025), prompting concern even as adoption continues Tiffin and Fraser (2025); Metnick et al. (2024). Further, the fragility of LLMs enables a range of misuses that could harm users Kumar and Lakkaraju (2024); Lin et al. (2025). For instance, Lin et al. (2025) showed that minor, unnoticed changes to suggested prompts can result in outputs with biases controlled by an adversary.
@@
Researchers have suggested that some of these issues correlate with the inconsistency of LLMs, which is generally defined as their tendency to generate low-confidence responses or conflicting responses when the same prompt is resampled Manakul et al. (2023). Accurately estimating LLM consistency is important because it can support multiple critical applications: predicting whether an answer is factual, thereby improving reliability in high-stakes domains Duan et al. (2024); detecting fragile or malicious prompts Lin et al. (2025); informing membership-inference attacks Mattern et al. (2023); or signaling to users the level of trust they should place in LLM outputs Kapoor et al. (2024); Ruggieri and Pugnana (2025).
@@
Consequently, many attempts have been made to define and measure the consistency of LLMs. This body of work can roughly be divided into two categories: (1) estimating consistency based on LLMs’ internal states or logits and (2) estimating consistency based on resampling LLMs’ responses. While the former is more computationally efficient, the latter is empirically well-grounded Kuhn et al. (2023); Qiu and Miikkulainen (2024) and applicable even without white-box access to a model Lin et al. (2024).
@@
Sampling-based estimation methods fundamentally rely on sampling from an LLM–prompt pair and comparing the outputs using a comparison function. This function differs between works—e.g., Duan et al. (2024) vs Manakul et al. (2023)—but partially boils down to estimating semantic similarity of responses. However, to our knowledge, none have based their estimation of LLM model consistency on user-based comparisons, the ground truth for semantic similarity Bowman et al. (2015); Agirre et al. (2014). Further, existing uncertainty estimation methods have not yet demonstrated reliability or alignment with human judgments of LLMs’ consistency. Without a baseline, consistency metrics are typically evaluated by their ability to predict whether a model output is factual Kuhn et al. (2023); Qiu and Miikkulainen (2024); Duan et al. (2024); Zhang et al. (2024); Kapoor et al. (2024). Joining recent work Novikova et al. (2025), we argue that consistency might be able to predict accuracy but is fundamentally an independent property of the model-prompt pair. There should be a metric that conveys to what degree a model, when provided a prompt, will repeatedly produce responses with equivalent meaning. We further posit that the baseline for measuring consistency is reliant on the comparison of semantics, the ground truth of which is defined by human judgement Bowman et al. (2015); Agirre et al. (2014); Nguyen et al. (2014); Raj et al. (2025).
@@
We aim to fill this gap by using user-based comparisons to estimate the consistency of LLMs, and by investigating whether an automated metric can come close to simulating users’ judgments. Specifically, we conduct a user study with 2,976 participants to collect semantic similarity ratings between a sample of 10 responses to each of 100 prompts, totaling 14,880 comparisons. We then calculate response-level (one score per response) and prompt-level (one score per prompt) consistency, establishing a user-based LLM consistency baseline. Through a series of experiments, we show that existing metrics of measuring LLM response uncertainty do not align well with human judgements collected in this study. We further show that an ensemble of logit-based scores is as similar to human judgement as the best-performing of the other methods we tested. We find that the discrepancy between existing metrics and human judgements fluctuates between models and between datasets, with previous metrics being less similar to human judgements on real-world prompts then on synthetic prompts. Based on our results, we advocate for more human-based LLM response consistency evaluation in future work.
@@
We structure the remainder of the paper in four sections: section˜2 details background work; section˜3 describes our model- and prompt-selection strategy, user study, and consistency calculation; section˜4 reports our experiments and comparisons; finally, we summarize and discuss implications of our work in section˜5.
@@ 2 Related Work
Here, we review work in consistency estimation of LLMs with resampling and internal model states. (2.1). We also cover work that investigated logit-based metrics to estimate LLM consistency (2.2). Finally, we reiterate our study motivation (2.3).
@@ 2.2 Probability- and Logit-Based Uncertainty Metrics
Existing work suggests that the confidence of individual token generation can suggest the consistency of LLM responses Manakul et al. (2023). Borrowing definitions from Manakul et al. (2023) , we denote the probability distribution of an individual token generated as $p$, and the entropy of individual token generation $H$, $H=-\sum_{i}p_{i}log(p_{i})$. Existing metrics include the average of the minus log probability Avg(log(p))Avg(-log(p)), maximum of the minus log probability Max(log(p))Max(-log(p)), average of entropy $Avg(H)$ and maximum of entropy $Max(H)$. In this paper, we ensemble these metrics (and variants) in 3.3.
@@
Besides probability-based (i.e., $p$-based) uncertainty metrics, we also find logit-based uncertainty metrics to suggest LLM response consistency. We denote the logits of individual token generation as $l$, $p=softmax(l)$. Specifically, we denote the largest, second largest, third largest, and fourth largest logit of a token generation as $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ correspondingly. The Difference of Logits Ratio Loss (DLR loss), formally $\frac{l_{1}-l_{2}}{[l_{1}-(l_{3}+l_{4})]/2}$, is a loss function commonly used by adversarial attacks to suggest how confident a model is Croce and Hein (2020). In this paper, we borrow the DLR loss as a logit-based uncertainty of a token generation to estimate the consistency of LLM responses.
Since probability-based uncertainty estimation is ultimately derived from token logits, for the rest of the paper, we call this group of uncertainty estimation logit-based methods.
@@ 2.3 Importance of User Evaluation in NLP Tasks
Human evaluations are important in improving and assessing the quality of natural language processing (NLP) Gritta et al. (2024); Boyd-Graber et al. (2022); Blodgett et al. (2024). Researchers have investigated how well machines are able to translate sentences into another language, compare sentences for semantic similarity, and many other NLP tasks with human evaluations Chatzikoumi (2020); Graham et al. (2017). This trend has continued in the LLM era Ouyang et al. (2022); Bai et al. (2022) While prior work has found ways to estimate the consistency of responses produced by LLMs, we argue that consistency is defined by what degree humans see responses as similar to each other. Using this definition, we establish a baseline through a user study.
@@ 3 Methods
In this study, we seek to understand how accurately existing metrics can approximate users’ perceptions of LLM consistency. We explore ways to create novel and more efficient ways of estimating LLM consistency without the need for expensive user studies or generating multiple responses per prompt. We provide an overview of our study design in figure˜1.
In this section, we first describe the prompts, models, and responses used in the study (3.1). We then describe in detail how we conducted the user study with 2,976 participants (3.2). Finally, we explain how we calculate consistency, both with existing metrics and our user-based method (3.3).
@@ 3.1 Prompts, Models, and Responses
Following prior work Geng et al. (2024), we selected two prompt sets: (1) to ensure comparability with prior research, we use a dataset of open-ended questions from CoQA Reddy et al. (2019), which is commonly adopted in the literature; and (2) to better represent real-world use cases, we include a sample of 50 prompts from LMSYS-Chat-1M Zheng et al. (2023).
Each entry in the CoQA dataset contains a story with a series of questions about the story. For our study, we only used the first question in each series, as many following questions are dependent on the previous question or answer for context. We randomly sampled 50 story–question pairs from the CoQA dataset.
Real-world prompts from the LMSYS dataset are more diverse than those in CoQA. Hence, we used labels in the data to filter out prompts that are were not in English or flagged111by OpenAI Moderation Zheng et al. (2023) (e.g., harassment, violence). Next, we randomly sampled 100 prompts from the filtered LMSYS dataset and manually removed 48 prompts: 14 coding questions, 9 about the LLM itself, 9 that were inappropriate (e.g., sexual), 6 that contain nonsensical sentences, 5 that we expected to have non-English responses (e.g., translations), 3 asking for answers over 1,500 words, and 2 that contain time-sensitive information. We excluded coding questions since participants might not have the background to rate coding-related responses. We randomly removed 2 prompts from the remaining 52 to obtain a final sample of 50.
@@
A generalizable definition of consistency should be model-agnostic. Thus, we used three open-weights models from competing institutions to generate responses: Llama-3.2-3B-Instruct (Llama), Gemma-2-9B-it (Gemma), and Mistral-7B-Instruct-v0.3 (Mistral). Each was the most recent publicly available version of their repective families, was widely used in previous work, and met our hardware memory constraints (Nvidia RTX A6000). We used open-weights models as they provided us access to logits, which were necessary for our calculation described in section˜3.3, and enabled reproducibility of our study Ma et al. (2024). We left temperature settings unchanged from the default configuration because there is no consensus from pervious work on optimal temperature for uncertainty estimation Cecere et al. (2025); Du et al. (2025); Wang et al. (2023); Renze and Guven (2024); Zou et al. (2023)
@@
The 100 prompts were randomly assigned to the three LLMs; Gemma received 34, Llama 33, and Mistral 33. Using the assigned model, we then generated 10 responses per prompt, totaling 1,000 responses. We took this approach as previous work showed that a set of 10 responses adequately captures the semantic diversity of the response space for uncertainty estimation purposes Kuhn et al. (2023); Qiu and Miikkulainen (2024). Response generation took 30 minutes of GPU time.
@@ 3.2 User Study
While many metrics have been created to approximate LLM consistency, a reporting of LLM consistency based on users’ perceptions has been largely missing from the literature. Our user-rating based approach establishes a baseline against which other metrics can be evaluated. We recruited users from Prolific and required that participants be 18 years or older, reside in the United States, read and type in English fluently, and have at least a 95% approval rate on the platform. We use Prolific over MTurk because recent work has shown its superior data quality Tang et al. (2022); Moss and Litman (2018).
Each participant was given an introduction to the study before seeing instructions for, and examples of, how to rate semantic similarity between a pair of sentences using a 6-point scale (Appendix A.1). We adopted this commonly-used scale, with explanations and examples, from Agirre et al. (2014). After the instructions, each participant rated the semantic similarity of five pairs of responses from five different prompts. Participants were compensated \$1; the median survey completion time was 5 minutes and 13 seconds. We collected participants’ Prolific IDs for compensating those who completed our survey. We removed Prolific IDs before doing any analysis. Lastly, we collected participants’ demographics. The sentence pairs and the order in which they were shown to participants were randomized. Our study was approved by our institution’s ethics review board.
@@
For each prompt, we generated 10 responses, yielding $\binom{10}{2}=45$ pairs of responses. With 100 prompts, we obtained a total of 4,500 pairs of responses, of which 2,938 were unique. More details about the number of prompts, responses, and pairs of unique responses for each model are described in table˜1.
@@ 3.3 Calculating and Comparing Consistency
Here, we outline how we calculated consistency scores (we calculated similarities between pairs of responses, consistency for each response, and consistency for each prompt) and compared user-based scores to previous work.
@@
We first aggregate user ratings of similarity between a pair of responses $r_{a}$ and $r_{b}$ by averaging ratings of $n\geq 5$ participants after removing the highest and lowest scores Curran (2016). Formally,
@@
To understand how existing metrics used for LLM consistency compare to human ratings, we compared them at the response-pair level. For each prompt (e.g., “What is Wh in batteries?”), 10 responses yielded 45 response-pairs, totaling 4,500 response-pairs for the 100 prompts in our study. For these response-pairs, we calculated the Spearman correlation coefficient $\rho$ between user ratings of similarity ($s_{\text{user}}$) and each of the four existing metrics: Bert ($s_{\text{\emph{Bert}}}$), BLEU ($s_{\text{\emph{BLEU}}}$), Rouge ($s_{\text{\emph{Rouge}}}$), and USE ($s_{\text{\emph{USE}}}$). We used the Spearman correlation coefficient $\rho$ because human ratings were on a 6-point Likert scale, and we did not assume linear correlation between human ratings and existing metrics. We share our findings on how each of these compare to human ratings in section˜4.2.
We calculated $s_{\text{\emph{Bert}}}$ using the BERTScore python package [^2] https://pypi.org/project/bert-score/0.3.0/, $s_{\text{\emph{BLEU}}}$ using the NLTK python package [^3] https://www.nltk.org/_modules/nltk/translate/bleu_score.html, $s_{\text{\emph{Rouge}}}$ using the rouge-score python package [^4] https://pypi.org/project/rouge-score/, and $s_{\text{\emph{USE}}}$ using the universal-sentence-encoder model on Kaggle [^5] https://www.kaggle.com/models/google/universal-sentence-encoder.
@@
We used the similarity scores between pairs of responses to calculate the consistency between a specific response $r_{i}$ and all the remaining m1m-1 responses. We call this response to response-set consistency, and compute it by averaging the response-to-response similarities between the specific response and each of the other responses. Formally,
@@
In addition to the four metrics used in response-pair level comparison, we add Semantic Entropy (SEKuhn et al. (2023) at this level of comparison, as their definition of consistency is measured at the response to response-set level. To understand how well these five metrics perform at this level, we compute the Spearman correlation coefficient and the mean squared error between each of the metrics and human ratings in section˜4.3.
Additionally, motivated by reducing the overhead of resampling, we compare logits-based scores to human ratings and propose an ensembling method that uses a linear combination of logits in section˜4.3.
To calculate SE, we used the roberta-large-mnli model to check entailment between pairs of responses. [^6] https://huggingface.co/FacebookAI/roberta-large-mnli We provide more details on logit-based scores and our ensembling method in subsection 3.3.
@@
Building on response to response-set consistency, we define prompt consistency as the average of the $m$ response to response-set consistency values of all responses to the prompt; formally,
@@
We use the Spearman correlation and mean squared error to determine how well each of the existing metrics and our ensemble approximate human ratings of prompt level consistency. We share our findings in section˜4.4.
Estimating consistency with logit-based metrics
We estimate the consistency of LLM responses using token-based uncertainty metrics used in previous work Manakul et al. (2023). Specifically, we use four uncertainty metrics: the probability, minus log probability, entropy, and the DLR loss (introduced in 2.2). For each response, we measure the maximum, sum, minimum, and average of these four token-based metrics, totaling $16$ values. The maximum and minimum show the ranges on individual tokens, while the average is the expected value across all tokens. The sum suggests a cumulative estimate over the whole response, as different responses may have different lengths. Unlike existing work that detected hallucination based producing a single response Manakul et al. (2023), we collect several responses to estimate LLM consistency for a prompt, corresponding to the definition in equation˜3.
@@
In addition to individual uncertainty-based metrics described in the previous paragraph, we create an ensemble of these in an attemp to better approximate human ratings of LLMs’ consistency. We used Sequential Feature Selection Rückstieß et al. (2011) to determine the most important metrics before composing them into an ensemble.